OpenStudy (anonymous):

f(x)=4x+7 and g(x)=3x-5 find (g)(f)(-4)

OpenStudy (e.mccormick):

So, how did you set it up?

OpenStudy (lynfran):

\[f(x)=4x+7, g(x)=3x-5 \]\[g(f)=3(4x+7)-5\]\[12x+21-5\]12x+16 \[g(f)(-4)=12(-4)+16=-32\]